# Cognitive Effort in the Two-Step Task: An Active Inference Drift-Diffusion Model Approach

### v1

@@ 3.4 Expected Free Energy discriminability affects second stage, but not first stage reaction times
To evaluate the models’ goodness-of-fit and predictive accuracy, we compared model-simulated behaviour on the two-step task to the observed behaviour of participants. This was done using a decile-binned analysis based on the absolute difference in the net EFE between the two actions, $|\Delta G^{\text{net}}|$. For each AIF-DDM model (except $\text{AIF-DDM}_{\text{NUD}}$), we first computed the trial-by-trial $|\Delta G^{\text{net}}|$ for each participant for both stages, based on their best-fit model parameters, choice history, and the experienced sequence of rewards and transitions. From the resulting distribution of $|\Delta G^{\text{net}}|$ values, for each participant and stage, we identified the decile boundaries ($10^{\text{th}}$-$90^{\text{th}}$ percentiles). We then binned each trial’s observed reaction time (RT) and choice into one of ten decile bins according to its $|\Delta G^{\text{net}}|$ value. For example, a trial with a $|\Delta G^{\text{net}}|$ value below the $10^{\text{th}}$ percentile was placed in the first bin ($0^{\text{th}}$-$10^{\text{th}}$). Within each bin, we calculated two metrics for each participant: (1) the mean RT, and (2) the probability of choosing the action with the lower net EFE, $P(\text{choose}\;G^{\text{net}}_{\text{min}})$. These participant-level metrics were then averaged across participants for each decile bin. Finally, for each trial stage, we generated 100 simulations to estimate the model’s predicted RT and choice probability. These simulated metrics were then binned and averaged in the same way as the observed data.
The simulation analysis results were qualitatively similar across all models, therefore we only display the predictions of the $\text{AIF-DDM}_{\text{FULL}}$ model as an example in Figure 3. Equivalent plots for the remaining models are presented in the Appendix 0.F. As shown in Figure 3 (top row), in both stages, participants were more likely to select the action with the minimum net EFE as the value of $|\Delta G^{\text{net}}|$ increased. This trend was well-captured by the AIF-DDM models.
The AIF-DDM models also predict that the mean RT should decrease as a function of EFE discriminability (i.e., as $|\Delta G^{\text{net}}|$ increases) for both stages. This effect can be observed for the second stage, however, it is almost negligible for the first one (Figure 3, bottom left).
@@
One potential explanation for the first-stage RT mismatch is the 2-second deadline. As mentioned in Section 2.1, all missed trials were excluded from the analysis, effectively truncating the observed RT distribution. In contrast, the AIF-DDM models lack this time restriction and thus predict a non-truncated distribution, which results in higher RT means (for further details, see Appendix 0.G). However, it is unlikely that this effect can fully account for the first-stage RT misfit because the second stage had the same deadline, yet a far less pronounced mismatch.
@@
3.5 The lack of EFE discriminability effect on first-stage reaction times might be explained by the experimental design
@@
The difference in inter-choice intervals could also explain why mean first-stage RTs are shorter than mean second-stage RTs (see Figure 3), which may seem counterintuitive since according to AIF, participants in the first stage should engage in planning, a more costly cognitive process than the one-shot second-stage decision.
@@ 4 Discussion
In this study, we introduced a novel AIF-DDM and evaluated its predictions in the two-step task. We discussed the difference between the model predictions and a recently developed information-theoretic formalism of cognitive effort based on AIF (Eq 4, [6]).
Theoretically, our model predicts that both the need to overcome a pre-existing habit and the difficulty of discriminating between options with similar subjective values (EFEs) increase cognitive effort during the first stage of the two-step task. We argue that for the first stage, both effects can be regarded as a single competition process between the ’habit-corrected’ or net EFEs of the available options. For the second stage, habits are assumed to play no role, and therefore, the model predicts that cognitive effort only depends on the discriminability between pure context-sensitive EFEs.
We found empirical evidence that EFE discriminability affects second-stage RTs of the two-step task as predicted by the AIF-DDM model - an effect that is not captured by the information-theoretic formalism. However, we found that the discriminability between the habit-corrected EFEs had a negligible effect on the first-stage RTs, contrary to the AIF-DDM predictions. We argue that the latter observation could be a result of the long inter-trial intervals that may allow participants to "pre-think" their choice before being presented with the options. Thus, first-stage RTs may only reflect motor execution, not the preceding cognitive effort.
@@
Our study has multiple limitations. Firstly, we focused on comparing the predictions of AIF-DDM with a single information-theoretic definition of cognitive effort. Our intention, however, is not to disprove this particular formalism, but to illustrate that a more complete theory of cognitive effort should include both the effect of subjective value competition and the violation of prior expectations. Secondly, we focused on studying how cognitive effort may result from the competition between the EFEs and prior beliefs of the different available options. However, we did not consider the cost of generating EFEs, which is likely higher during the multi-step planning required at the first stage compared to the one-shot second-stage choice. Finally, cognitive effort is a complex phenomenon that may depend upon multiple factors, such as the objective cognitive demand of a task or the participant’s information-processing speed. Some of these factors may be captured by different AIF-DDM parameters (e.g., $a_{\text{bs}}$ and $\kappa$ for impulsiveness and motivation, respectively); however, this analysis was beyond the scope of this paper.
@@ Appendix 0.A Further details on Active Inference in the two-step task
According to the learning rule implemented by [3], after observing the outcome of a chosen action, its corresponding $\alpha$ and $\beta$ parameters are updated following this rule:
@@
Where $\chi_{t}\in[0,1]$ is the surprise-modulated adaptation rate for trial $t$ which depends on the predictive surprise $PS=-\ln{p(o_{t};\pi_{t}(\theta))}$, and the prior volatility parameter $\nu_{\text{PS}}\in[0,1]$ which modulates the influence of surprise on the agent’s beliefs. In Eqs. 8 and 9, $l$ is the learning rate parameter and $\nu_{\text{SD}}$ is the volatility parameter of chosen actions and $\nu_{\text{SD}}$ modulates how over time $\alpha$ and $\beta$ decay towards its prior values $\alpha_{0}$ and $\beta_{0}$.
Evidence from multi-armed bandit tasks suggests a bias in prior outcome probabilities. To model this effect, [3] fit a free parameter to each participant $p_{r}$, which determine the initial (or uninformed prior) $\alpha$ and $\beta$ values, such that $\alpha_{0}=2(1-p_{r})$ and $\beta_{0}=2p_{r}$.
The decay effect of the chosen actions captured by $\nu_{\text{SD}}$, may also affect unchosen actions, and therefore, for these, the updating rule reduces to:
@@
The surprise update mechanism for chosen actions and the decay mechanisms for chosen and unchosen actions may not be all simultaneously implemented [3]. Therefore, four variants of the AIF models can be implemented, each with a different combination of update rules as pointed out in Section 2.2.
@@ Appendix 0.B Hybrid Reinforcement Learning Drift-Diffusion model
$Q^{\text{MF}}_{t+1}(s_{2,t},a_{2,t})=Q^{\text{MF}}_{t}(s_{2,t},a_{2,t})+\alpha_{2}\delta_{2,t}$ (13)
Where $\alpha_{2}$ is the second-stage learning rate parameter ($0\leq\alpha_{2}\leq 1$) and $\delta_{2,t}$, is the second-stage prediction error, defined as the difference between the observed outcome and its state-action value, such that $\delta_{2,t}=o_{t}-Q^{\text{MF}}_{t}(s_{2,t},a_{2,t})$
@@
$Q^{\text{MF}}_{t+1}(s_{\text{A}},a_{1,t})=Q^{\text{MF}}_{t}(s_{\text{A}},a_{1,t})+\alpha_{1}[\delta_{2,t}+\lambda\delta_{1,t}]$ (14)
Where $\alpha_{1}$ is the first-stage learning rate parameter ($0\leq\alpha_{1}\leq 1$), $\delta_{1,t}$, is the first-stage prediction error, defined as $\delta_{1,t}=Q^{\text{MF}}_{t}(s_{2,t},a_{2,t})-Q^{\text{MF}}_{t}(s_{\text{A}},a_{1,t})$, and $\lambda$ is the eligibility parameter ($0\leq\lambda\leq 1$), which determines how much the second-stage prediction error influences the first-stage state-action value.
For the model-based strategy, agents are assumed to have a generative model of the task, which in this case means that they know (or learn) the state transition probabilities and use these to calculate state-action values. Since the outcome of the second-stage actions does not depend. However, for the first-stage actions, the model-based values will be given by the following equation:
$Q_{t}^{\text{MB}}(s_{\text{A}},a_{1,t})=\sum_{s_{2}\in\{s_{\text{B}},s_{\text{c}}\}}P(s_{2}|s_{\text{A}},a_{1,t})\max_{a_{2}\in\mathcal{A}}Q_{t}^{\text{MB}}(s_{2},a_{2})$ (15)
@@
Where $w\in[0,1]$ is a weight parameter that regulates the relative influence of model-based and model-free strategies in the decision-making. The last term in the equation is added to model the tendency observed in participants to repeat the first-stage choice of the previous trial. The level of response stickiness is regulated by the parameter $\rho$, which multiplies the $\text{rep}(a)$, which is 1 if the first-stage action is the same as the one chosen in the previous trial and 0 otherwise.
At the second stage, the state-action values are computed using only a model-free strategy, and it is assumed that there is no tendency to repeat the previous action. Therefore, for the second stage, the net state-action value reduces to $Q^{\text{net}}_{t}(s_{2},a_{2})=Q^{\text{MF}}_{t}(s_{2},a_{2})$ the second-stage choice probabilities are given by the following distribution:
Once state-action values are computed, the HRL-DDM agent, will pick an action $a$ at trial $t$, state $s$ and stage $p=\{1,2\}$ in a drift-diffusion process with a drift rate given by the following equation:
@@
Where $\mathcal{A}_{s}$ are these of actions available at state $s$
@@ Appendix 0.C Further details on the model fitting and comparison procedures
For the two-step task, the MLE procedure consists of finding the participant’s set of parameter values of the model $m$, $\hat{\theta}_{m}^{\mathrm{MLE}}$ that maximizes the likelihood of the first-stage $d^{1}_{1:T}$ and second-stage $d^{2}_{1:T}$ behavioural data, given the parameters set $p(d^{1}_{1:T},d^{2}_{1:T}|\theta_{m},m)$, or equivalently, Log-Likelihood, $LL=-\log\,p(d_{1:T}|\theta_{m},m)$. This quantity can be expressed as the sum of the individual trial LLs [17], such that:
@@
For AIF and HRL models, $d^{p}_{t}$, corresponds to the choice made at stage $p$ and trial $t$, so Eq. 19 can be evaluated using the corresponding choice probability distributions of each model. For AIF-DDM and HRL-DDM, however, $d^{p}_{t}$, corresponds to the choice and RT tuple of stage $p$ and trial $t$. Hence, compute the Likelihood of each trial and phase using other Wiener’s First-Passage Distribution using the HDDM Python package. The WFPT distribution takes as input the parameters $a_{bs}$, $t_{nd}$ well as the drift rates, which for HRL-DDM and AIF-DDM are calculated using Equations 17 and 3, respectively, which are calculated using the rest of the free parameters.
@@ Appendix 0.E Parameter recovery results
(a) $\text{AIF(-DDM)}_{\text{FULL}}$
@@ Appendix 0.G Reaction Time Q-Q Analysis
To further investigate the reaction time (RT) misfit mentioned in Section 3.4, we generated Quantile-Quantile (Q-Q) plots comparing observed versus simulated within-participant RT deciles ($10^{th}-90^{th}$). Simulated RT deciles were produced for each participant by running 50 simulated experiments with 1000 trials each (with the same reward volatility as the real experiment) using the best-fitted parameters. As shown in Figure 7, the plots reveal that all models consistently overestimate the higher RT quantiles for both stages. This overestimation was expected, as the models’ predictions are unconstrained, whereas the participants’ observed RTs were truncated by a 2-second response deadline, with missed trials excluded from the analysis.
Figure 7: Systematic overestimation of higher reaction time (RT) quantiles by the AIF-DDM models. The Q-Q plots compare observed versus predicted RT deciles for Stage 1 (left) and Stage 2 (right). Points lying above the identity line (red) indicate that the model’s predicted RT is higher than the observed RT. This effect is most pronounced for the longest reaction times, where predictions deviate furthest from the identity line. Note that the purple markers indicate an overlap of red and blue points.

### v2

@@ 3.4 Expected free energy discriminability affects second stage, but not first stage reaction times
To evaluate the models’ goodness-of-fit and predictive accuracy, we compared model-simulated behaviour on the two-step task to the observed behaviour of participants. This was done using a decile-binned analysis based on the absolute difference in the net EFE between the two actions, $|\Delta G^{\text{net}}|$. For each AIF-DDM model (except $\text{AIF-DDM}_{\text{NUD}}$), we first computed the trial-by-trial $|\Delta G^{\text{net}}|$ for each participant for both stages, based on their best-fit model parameters, choice history, and the experienced sequence of rewards and transitions. From the resulting distribution of $|\Delta G^{\text{net}}|$ values, for each participant and stage, we identified the decile boundaries ($10^{\text{th}}$-$90^{\text{th}}$ percentiles). We then binned each trial’s observed RT and choice into one of ten decile bins according to its $|\Delta G^{\text{net}}|$ value. For example, a trial with a $|\Delta G^{\text{net}}|$ value below the $10^{\text{th}}$ percentile was placed in the first bin ($0^{\text{th}}$-$10^{\text{th}}$). Within each bin, we calculated two metrics for each participant: (1) the mean RT, and (2) the probability of choosing the action with the lower net EFE, $P(\text{choose}\;G^{\text{net}}_{\text{min}})$. These participant-level metrics were then averaged across participants for each decile bin. Finally, for each trial stage, we generated 100 simulations to estimate the model’s predicted RT and choice probability. These simulated metrics were then binned and averaged in the same way as the observed data.
The simulation analysis results were qualitatively similar across all models, therefore we only display the predictions of the $\text{AIF-DDM}_{\text{FULL}}$ model as an example in Fig. 3. Equivalent plots for the remaining models are presented in the Appendix 0.F. In both stages, participants were more likely to select the action with the minimum net EFE as the value of $|\Delta G^{\text{net}}|$ increased. This trend was well-captured by all the AIF-DDM models (see Appendix 0.F).
The AIF-DDM models also predict that the mean RT should decrease as a function of EFE discriminability (i.e., as $|\Delta G^{\text{net}}|$ increases) for both stages. This effect can be observed for the second stage; however, it is almost negligible for the first one (Fig. 3, bottom left).
@@
Multiple explanations may account for the mismatch in first-stage RTs. The first concerns the 2-second deadline. As described in Section 2.1, all missed trials were excluded from the analysis, effectively truncating the observed RT distribution. In contrast, the AIF-DDM models do not impose this time restriction, resulting in higher predicted RT means (see Appendix 0.G for further details). However, this effect alone is unlikely to fully explain the misfit, since the second stage had the same deadline, yet a far less pronounced mismatch.
[…]
@@
3.5 The lack of net EFE discriminability effect on first-stage reaction times might be explained by the experimental design
@@
The difference in inter-choice intervals could also explain why mean first-stage RTs are shorter than mean second-stage RTs (see Fig. 3), which may seem counterintuitive since, according to AIF, participants in the first stage should engage in planning, a more costly cognitive process than the one-shot second-stage decision.
@@ 4 Discussion
In this study, we introduced a novel AIF-DDM and evaluated its predictions in the two-step task. We discussed the difference between the model predictions and a recently developed information-theoretic formalism of cognitive effort based on AIF (Eq. 4, [10]).
Our model predicts that both the need to overcome a pre-existing habit and the difficulty of discriminating between options with similar subjective values (EFEs) increase cognitive effort during the first stage of the two-step task. We argue that for the first stage, both effects can be regarded as a single competition process between the ’habit-corrected’ or net EFEs of the available options. For the second stage, habits are assumed to play no role, and therefore, the model predicts that cognitive effort only depends on the discriminability between pure context-sensitive EFEs.
We found empirical evidence that EFE discriminability affects second-stage RTs of the two-step task as predicted by the AIF-DDM model – an effect that is not captured by the information-theoretic formalism. However, we found that the discriminability between the habit-corrected EFEs had a negligible effect on the first-stage RTs, contrary to the AIF-DDM predictions. We argue that the latter observation could be a result of the long inter-trial intervals that may allow participants to "pre-think" their choice before being presented with the options. Thus, first-stage RTs may only reflect motor execution, not the preceding cognitive effort.
@@
Our study has multiple limitations. Firstly, we focused on comparing the predictions of AIF-DDM with a single information-theoretic definition of cognitive effort. Our intention, however, is not to disprove this particular formalism, but to illustrate that a more complete theory of cognitive effort should include both the effect of subjective value competition and the violation of prior expectations. Secondly, we focused on studying how cognitive effort may result from the competition between the EFEs and prior beliefs of the different available options. We did not, however, address the costs associated with learning or with generating EFEs, which are likely to differ between the two stages. Finally, cognitive effort is a complex phenomenon that may depend upon multiple factors, such as the objective cognitive demand of a task or the participant’s information-processing speed. Some of these factors may be captured by different AIF-DDM parameters (e.g., $a_{\text{bs}}$ and $\kappa$ for impulsiveness and motivation, respectively); however, this analysis was beyond the scope of this paper.
@@ Appendix 0.A Further details on Active Inference in the two-step task
According to the learning rule implemented by [4], after observing the outcome of a chosen action, its corresponding $\alpha$ and $\beta$ parameters are updated following these rules:
@@
Where $\chi_{t}\in[0,1]$ is the surprise-modulated adaptation rate for trial $t$ which depends on the predictive surprise $PS=-\ln{p(o_{t};\pi_{t}(\theta))}$, and the prior volatility parameter $\nu_{\text{PS}}\in[0,1]$ which modulates the influence of surprise on the agent’s beliefs. In Eqs. 8 and 9, $l$ is the learning rate parameter, $\nu_{\text{SD}}$ is the volatility parameter of chosen actions and $\nu_{\text{SD}}$ modulates how over time $\alpha$ and $\beta$ decay towards their prior values $\alpha_{0}$ and $\beta_{0}$, respectively.
Evidence from multi-armed bandit tasks suggests a bias in prior outcome probabilities. To model this effect, a free parameter, $p_{r}$, is fitted to each participant, which determines the initial (or uninformed) $\alpha$ and $\beta$ prior values, such that $\alpha_{0}=2(1-p_{r})$ and $\beta_{0}=2p_{r}$ [4].
The decay effect of the chosen actions captured by $\nu_{\text{SD}}$, may also affect unchosen actions, and therefore, for these, the updating rules reduce to:
@@
The surprise update mechanism for chosen actions and the decay mechanisms for chosen and unchosen actions may not all be simultaneously operating [4]. Therefore, four variants of the AIF models can be implemented, each with a different combination of update rules as pointed out in Section 2.2.
@@ Appendix 0.B Hybrid reinforcement learning drift-diffusion model
$Q^{\text{MF}}_{t+1}(s_{2},a_{2})=Q^{\text{MF}}_{t}(s_{2},a_{2})+\alpha_{2}\delta_{2,t}$ (13)
Where $\alpha_{2}\in[0,1]$ is the second-stage learning rate parameter and $\delta_{2,t}$, is the second-stage prediction error, defined as the difference between the observed outcome and its state-action value, such that $\delta_{2,t}=o_{t}-Q^{\text{MF}}_{t}(s_{2},a_{2})$
@@
$Q^{\text{MF}}_{t+1}(s_{\text{A}},a_{1})=Q^{\text{MF}}_{t}(s_{\text{A}},a_{1})+\alpha_{1}[\delta_{2,t}+\lambda\delta_{1,t}]$ (14)
Where $\alpha_{1}\in[0,1]$ is the first-stage learning rate parameter, $\delta_{1,t}$, is the first-stage prediction error, defined as $\delta_{1,t}=Q^{\text{MF}}_{t}(s_{2},a_{2})-Q^{\text{MF}}_{t}(s_{\text{A}},a_{1})$, and $\lambda\in[0,1]$ is the eligibility parameter, which determines how much the second-stage prediction error influences the first-stage state-action value.
For the model-based strategy, agents are assumed to have a generative model of the task, which in this case means that they know (or learn) the state transition probabilities and use these to calculate state-action values. For the first-stage actions, the model-based values depend on the second-stage model-free values, such that:
$Q_{t}^{\text{MB}}(s_{\text{A}},a_{1})=\sum_{s_{2}\in\{s_{\text{B}},s_{\text{C}}\}}P(s_{2}|s_{\text{A}},a_{1})\max_{a_{2}\in\mathcal{A}_{s_{2}}}Q_{t}^{\text{MF}}(s_{2},a_{2})$ (15)
@@
Where $w\in[0,1]$ is a weight parameter that regulates the relative influence of model-based and model-free strategies in the decision-making. The last term in the equation is added to model the tendency observed in participants to repeat the first-stage choice of the previous trial. The level of response stickiness is regulated by the parameter $\rho$, which multiplies the function $\text{rep}(a)$ – equal to 1 if the first-stage action is the same as the one chosen in the previous trial and 0 otherwise.
At the second stage, the state-action values are computed using only a model-free strategy, and it is assumed that there is no tendency to repeat the previous action. Therefore, for the second stage, the net state-action value reduces to $Q^{\text{net}}_{t}(s_{2},a_{2})=Q^{\text{MF}}_{t}(s_{2},a_{2})$.
Once state-action values are computed, the HRL-DDM agent, will pick an action $a$ at trial $t$, state $s$ and stage $p=\{1,2\}$ through a drift-diffusion process with a drift rate given by the following equation:
@@
Where $\mathcal{A}_{s}$ is the set of available actions at state $s$.
@@ Appendix 0.C Further details on the model fitting and comparison procedures
For the two-step task, the MLE procedure consists of finding the participant’s set of parameter values of the model $m$, $\hat{\theta}_{m}^{\mathrm{MLE}}$, that maximizes the likelihood of the first-stage and second-stage behavioural data ($d^{1}_{1:T}$ and $d^{2}_{1:T}$, respectively), given the parameters set, $p(d^{1}_{1:T},d^{2}_{1:T}|\theta_{m},m)$, or equivalently, log-likelihood, $LL=\log\,p(d^{1}_{1:T},d^{2}_{1:T}|\theta_{m},m)$. This quantity can be expressed as the sum of the individual trial LLs [23], such that:
@@
For AIF and HRL models, $d^{p}_{t}$, corresponds to the choice made at stage $p$ and trial $t$, so Eq. 19 can be evaluated using the corresponding choice probability distributions of each model. For AIF-DDM and HRL-DDM, however, $d^{p}_{t}$, corresponds to the choice and RT tuple of stage $p$ and trial $t$. Hence, to compute the likelihood of each trial and phase, we use the WFPT distribution as implemented by the HDDM Python package. The WFPT distribution takes as input the parameters $a_{bs}$, $t_{nd}$, as well as the drift rates, which for HRL-DDM and AIF-DDM are calculated using Eqs. 17 and 3, respectively.
@@ Appendix 0.E Parameter recovery results
((a)) $\text{AIF(-DDM)}_{\text{FULL}}$
@@ Appendix 0.G Reaction time Q-Q analysis
To further investigate the reaction time (RT) misfit mentioned in Section 3.4, we generated Quantile-Quantile (Q-Q) plots comparing observed versus simulated within-participant RT deciles ($10^{th}-90^{th}$). Simulated RT deciles were produced for each participant by running 50 simulated experiments with 1000 trials each (with the same reward volatility as the real experiment) using the best-fitted parameters. As shown in Fig. 7, the plots reveal that all models consistently overestimate the higher RT quantiles for both stages. This overestimation was expected, as the models’ predictions are unconstrained, whereas the participants’ observed RTs were truncated by a 2-second response deadline, with missed trials excluded from the analysis.
Figure 7: Systematic overestimation of higher reaction time (RT) deciles by the AIF-DDM models. The Q-Q plots compare observed versus predicted RT deciles for Stage 1 (left) and Stage 2 (right). Points lying above the identity line (red) indicate that the model’s predicted RT is higher than the observed RT. This effect is most pronounced for the longest reaction times, where predictions deviate furthest from the identity line. Note that the purple markers indicate an overlap of red and blue points.